Trina Appleton, Jackson's Food Marts marketing director, read from a national grocery magazine that the average wait time at a convenience food mart is 2.5 minutes. She took a random sample of 20 shoppers at the Jackson's Food Mart on Overland Road. The average wait for the sample shoppers was 2.8 minutes with a standard deviation of 1.5 minutes. She asked you if the average wait time at the Overland Jackson's was really less than the national average stated in the magazine or was it just sampling error. Run a hypothesis test on the sample with a 0.05 significance level and write a memo the Ms. Appleton stating your conclusion. (Hint: This is a lower-tailed, t-test.)
Answer #1

Sol:

H0:μ = 2.5

Ha:mu <2.5

alpha=0.05

test statistic

t=xbar-mu/s/sqrt(n)

=(2.8-2.5)/(1.5/sqrt(20))

=0.89

t=0.893

df=n-1=20-1=19

p=1-T.DIST.RT(0.893;19)=0.8087

p>0.05

Reject H0.
Accept Ha

Conclusion:

there is sufficient statistical evidence at 5% level of significance to conclude that

the average wait time at the Overland Jackson's was really less than the national average stated in the magazine